# Learning-Based Modeling of a Magnetically Steerable Soft Suction Device for Endoscopic Endonasal Interventions

### v2

@@ 4.3 Feature Importance Analysis
To gain insights into the relationship between magnetic field parameters and device deformation, we analyzed the feature importance metrics derived from the RF model. Our results revealed distinct patterns of influence across different control points: The magnetic field components ($b_{x}$ and $b_{y}$) dominate the influence on most control point coordinates, with $b_{y}$ showing particularly high importance (>>50%) for ${p}_{0Y}$, ${p}_{2Y}$, and ${p}_{3Y}$ coordinates. This aligns with the physical understanding of magnetic actuation, where the Y-component of the magnetic field primarily controls lateral movement of the device. The frequency parameter exhibits uniquely high importance (32.34%) for ${p}_{0Z}$, suggesting that actuation frequency plays a significant role in determining the base height of the device. The distance parameter shows consistent influence on $\mathbf{p}_{0}$ coordinates (23.81% for ${p}_{0X}$, 24.50% for ${p}_{0Z}$) and $\mathbf{p}_{3}$ coordinates (13.00% for ${p}_{3X}$, 13.17% for ${p}_{3Y}$, 15.79% for ${p}_{3Z}$), reflecting how the proximity to the electromagnetic coil system affects both the fixed base and the free tip of the device. This feature importance analysis provides valuable insights for control system design, indicating which magnetic field parameters should be prioritized for precise manipulation of specific device coordinates. For instance, fine control of the device tip position would benefit from careful modulation of both $b_{x}$ and $b_{y}$, while actuation frequency could be leveraged for height adjustment at the device base.
@@ 4.4 Bezier Curve Shape Reconstruction
Figure 11: Bezier curve shape prediction (3D view). Ground truth (blue) vs. RF predictions (green).
@@ 4.5 FBG Ground Truth Limitations and Accuracy
The accuracy validation methodology of this study warrants careful examination, particularly regarding the use of reconstructed shapes from FBG as ground truth for training the learning based models. While multi-core fiber Bragg grating sensors are well-established for shape sensing in medical robotics applications, their inherent limitations must be acknowledged when interpreting the reported accuracy metrics. The FBG system employed in this study is notably coarser than comparable studies, such as Dong et al. [33], who achieved 3.17 mm grating spacing and validated their FBG measurements against electromagnetic tracking, reporting 1.53 mm tip position accuracy. The piecewise constant curvature assumption between grating points introduces interpolation errors that accumulate along the fiber length, potentially resulting in tip position uncertainties exceeding 1 mm despite the sensor’s high strain sensitivity.
The reported sub-millimeter accuracy (0.064 mm RMSE) in this work specifically quantifies the Random Forest model’s ability to predict Bezier control points that, when used to reconstruct the shape, closely match the FBG-derived measurements. This metric should not be interpreted as absolute shape sensing accuracy, but rather as the model’s capacity to learn and reproduce the mapping between magnetic field inputs and FBG-measured deformations. The distinction is critical: while the learning-based approach successfully captures the nonlinear relationship between actuation and measured response with high fidelity, the absolute accuracy of the predicted shapes relative to true physical deformation remains bounded by the FBG system’s inherent limitations. Future work should incorporate independent validation using calibrated vision systems or higher resolution tracking methods to establish absolute accuracy bounds. Despite these limitations, the learning-based approach remains valid for the intended application of closed-loop control, where consistent relative shape prediction is more critical than absolute position accuracy. The model’s ability to predict deformation patterns with high repeatability enables effective navigation control even if absolute position uncertainty exists, as the controller can compensate for systematic errors through feedback mechanisms.
@@ 5 CONCLUSION
This paper presented the first learning-based approach that transforms FBG shape sensing data into Bezier control points for modeling magnetically actuated soft surgical instruments. This novel approach bridges the gap between shape sensing technologies and practical control applications by providing a direct mapping between the magnetic field and the resulting device configuration. Our key contribution was the development and validation of a framework that enables highly accurate real-time shape prediction of complex non-linear deformations in soft robotic devices. By comparing NN and RF modeling approaches, we demonstrated that ensemble-based methods outperform traditional NN for this application, particularly at the critical distal end of the device, where precise control is essential for surgical tasks. Also, the computational efficiency of our approach supports high-frequency real-time control suitable for dynamic surgical environments, while maintaining sub-millimeter accuracy across the device’s entire structure. Future work will focus on three main directions: (1) integrating this modeling approach into a closed-loop control system for autonomous navigation, (2) expanding the training dataset to include additional device configurations for enhanced generalization, and (3) validating the model through ex vivo experiments in simulated surgical scenarios.

### v3

@@ 3.3 Feature Importance Analysis
To gain insights into the relationship between magnetic field parameters and device deformation, we analyzed the feature importance metrics derived from the RF model. Our results revealed distinct patterns of influence across different control points: The magnetic field components ($b_{x}$ and $b_{y}$) dominate the influence on most control point coordinates, with $b_{y}$ showing particularly high importance (>>50%) for ${p}_{0Y}$, ${p}_{2Y}$, and ${p}_{3Y}$ coordinates. This aligns with the physical understanding of magnetic actuation, where the Y-component of the magnetic field primarily controls lateral movement of the device. The frequency parameter exhibits uniquely high importance (32.34%) for ${p}_{0Z}$, suggesting that actuation frequency influences the vertical position (z-coordinate) of the device’s proximal segment, i.e., the slight vertical displacement at the base resulting from dynamic bending under oscillating magnetic fields. The distance parameter shows consistent influence on $\mathbf{p}_{0}$ coordinates (23.81% for ${p}_{0X}$, 24.50% for ${p}_{0Z}$) and $\mathbf{p}_{3}$ coordinates (13.00% for ${p}_{3X}$, 13.17% for ${p}_{3Y}$, 15.79% for ${p}_{3Z}$), reflecting how the proximity to the electromagnetic coil system affects both the fixed base and the free tip of the device. This feature importance analysis provides valuable insights for control system design, indicating which magnetic field parameters should be prioritized for precise manipulation of specific device coordinates. For instance, fine control of the device tip position would benefit from careful modulation of both $b_{x}$ and $b_{y}$, while actuation frequency could be leveraged for height adjustment at the device base.
@@ 3.4 Bezier Curve Shape Reconstruction
Figure 11: Bezier curve shape prediction (3D view). Ground truth shown in blue and RF predictions shown in red. The close alignment between the predicted and measured curves demonstrates the RF model’s strong capability to capture complex nonlinear deformations of the soft suction device. The figure highlights the base and tip of the robot, illustrating how the predicted Bezier curve accurately preserves both overall curvature and tip position within sub-millimeter deviation.
@@ 3.5 FBG Ground Truth Limitations and Accuracy
The accuracy validation methodology of this study warrants careful examination, particularly regarding the use of reconstructed shapes from FBG as ground truth for training the learning based models. While multi-core fiber Bragg grating sensors are well-established for shape sensing in medical robotics applications, their inherent limitations must be acknowledged when interpreting the reported accuracy metrics. The FBG system employed in this study uses a multi-core fiber containing 26 gratings uniformly spaced at 10 mm intervals along a total sensing length of 250 mm, with approximately five gratings located within the 40 mm segment embedded in the soft body of the device. This configuration provides sufficient spatial resolution for curvature reconstruction. To provide context regarding sensor resolution capabilities in FBG-based shape sensing, we note that Dong et al. [10] employed a denser grating configuration (3.17 mm spacing) and reported 1.53 mm tip position accuracy when validated against electromagnetic tracking. While their experimental setup, validation methodology, and application context differ from the present study, this reference illustrates that finer grating resolution combined with independent external validation can achieve higher absolute positioning accuracy. The sensing layout used here is comparatively coarser, which motivated our decision to report model-to-FBG consistency rather than absolute shape accuracy. The piecewise constant curvature assumption between grating points introduces interpolation errors that accumulate along the fiber length, potentially resulting in tip position uncertainties exceeding 1 mm despite the sensor’s high strain sensitivity.
The reported sub-millimeter accuracy (0.064 mm RMSE) quantifies the fidelity of the Bezier curve representation relative to the FBG-reconstructed shape, confirming that the four-point parameterization accurately captures the device’s geometry. The RF model’s prediction accuracy for these control points, reported in Table 3, achieved an overall RMSE of 0.087 mm. The RMSE used throughout this study was computed as:
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left\|\mathbf{y}_{i}-\hat{\mathbf{y}}_{i}\right\|^{2}}$ (10)
where $\mathbf{y}_{i}$ and $\hat{\mathbf{y}}_{i}$ denote the ground-truth and predicted 3D point coordinates, respectively, and $N$ is the total number of evaluated points. This metric should not be interpreted as absolute shape sensing accuracy, but rather as the model’s capacity to learn and reproduce the mapping between magnetic field inputs and FBG-measured deformations. The distinction is critical: while the learning-based approach successfully captures the nonlinear relationship between actuation and measured response with high fidelity, the absolute accuracy of the predicted shapes relative to true physical deformation remains bounded by the FBG system’s inherent limitations. It should be noted that FBG shape reconstruction employs continuous shape reconstruction, implemented via piecewise constant curvature (PCC) interpolation between discrete grating locations, which is standard practice in FBG-based sensing systems. This is fundamentally different from traditional PCC-based kinematic modeling, where constant-curvature assumptions are used to analytically derive actuation-to-shape relationships. Our learning framework captures experimentally observed behaviors, including frequency-dependent and viscoelastic effects, that analytical PCC kinematic models cannot represent.
This distinction is further evidenced by our experimental results, where the same magnetic field magnitude produces different deformation patterns at different actuation frequencies (as shown in Fig. 6), a dynamic behavior that the learning model successfully captures but that analytical PCC kinematic models fundamentally cannot represent. The learning model thus encodes richer behavioral information than what analytical PCC-based models could provide, even though both share a common continuous shape reconstruction step.
Future work will incorporate independent validation using calibrated vision systems or higher-resolution tracking methods to establish absolute accuracy bounds. Despite these limitations, the learning-based approach remains valid for the intended application of real-time shape estimation, where consistent relative deformation tracking is more critical than absolute position accuracy. The model’s ability to predict deformation patterns with high repeatability demonstrates strong potential for integration into future closed-loop control frameworks, where feedback from embedded FBG sensing could be leveraged to compensate for systematic errors and improve precision during surgical manipulation. Additionally, future work could explore hybrid modeling strategies that combine the theoretical rigor of physics-based frameworks such as the GVS approach [26] with data-driven components for parameter identification or model correction, potentially offering improved generalizability while retaining the empirical accuracy demonstrated in this study.
@@
Also, the rotating magnetic field profiles employed in this study were selected to provide controlled and repeatable deformation trajectories while sampling different magnetic field directions across the coil workspace. As the tip moves through circular paths, the local magnetic field vectors vary spatially around the coil table, enabling the model to learn deformation responses under multiple actuation orientations. Additionally, the experiments were conducted across a frequency range of 0.2–1.0 Hz to capture the dynamic, viscoelastic behavior of the soft body under time-varying loads. This combined spatial and temporal variation provided a rich dataset for model training and validation. Future work will extend this framework to include non-circular and task-oriented magnetic actuation profiles, such as targeted tip positioning and complex navigation trajectories, to further evaluate the model’s generalization in conditions that closely resemble surgical operations.
@@
3.6 Generalization to confined surgical environments
The experiments presented in this work were performed in an open-space setting to isolate and characterize the fundamental deformation behavior of the magnetically actuated soft suction device. This environment provided a controlled platform for capturing the intrinsic relationship between magnetic actuation inputs and device deformation without external constraints, ensuring consistent data for model training and validation. However, in realistic endoscopic endonasal surgery, the device will operate within narrow anatomical corridors surrounded by rigid and soft tissues. These boundaries can restrict bending, cause asymmetric deformation, and introduce frictional or damping effects due to tissue contact and mucus layers. We acknowledge that such confinement and viscous interaction could alter the deformation patterns observed under open-space conditions. Therefore, the current results represent a baseline characterization of the device’s intrinsic magnetic response, which serves as a necessary first step before introducing the additional complexities of tissue interaction. To extend the modeling framework toward clinically relevant environments, future work will incorporate training data acquired within anatomically realistic phantoms that replicate skull-base geometries and the mechanical properties of nasal tissues. The existing Bezier-based learning model is compatible with such data expansion because it encodes shape using a compact set of control points that can adapt to new boundary conditions. During these experiments, additional input features such as contact location, frictional interaction indicators, or limited endoscopic vision feedback may be included to improve model robustness under partial confinement. Furthermore, the integration of real-time FBG sensing with image-based feedback could enable adaptive online correction of the predicted shape during surgical manipulation. Through these planned steps, the proposed learning-based framework can be systematically generalized to accurately predict and control device behavior in the constrained and viscoelastic environments encountered in actual endonasal neurosurgery.
@@ 4 Conclusion
This paper presented a generalized learning-based modeling framework that transforms FBG shape sensing data into Bezier control points, establishing a new abstraction for modeling magnetically actuated soft surgical instruments. This framework introduces a unified methodology that links magnetic actuation inputs, embedded sensing, and Bezier-based geometric representation into an interpretable and compact modeling structure. This approach bridges the gap between shape sensing technologies and practical control applications by providing a direct mapping between the magnetic field and the resulting device configuration. Our key contribution was the development and validation of a framework that enables highly accurate real-time shape prediction of complex non-linear deformations in soft robotic devices. By comparing NN and RF modeling approaches, we demonstrated that ensemble-based methods outperform traditional NN for this application, particularly at the critical distal end of the device, where precise control is essential for surgical tasks. Also, the computational efficiency of our approach supports high-frequency real-time control suitable for dynamic surgical environments, while maintaining sub-millimeter accuracy across the device’s entire structure. This study establishes a new modeling paradigm that integrates magnetic actuation, data-driven learning, and geometric Bezier parameterization, forming a transferable foundation for future control and design of magnetically actuated continuum robots. Future work will focus on three main directions: (1) integrating this modeling approach into a closed-loop control system for autonomous navigation, (2) expanding the training dataset to include additional device configurations for enhanced generalization, including exploration of hybrid physics-informed learning approaches that leverage established frameworks such as GVS for improved interpretability and generalizability, and (3) validating the model through ex vivo experiments in simulated surgical scenarios.